OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (0, 5) and (–2, 11)?

OpenStudy (anonymous):

first find the slope, then using one of your given points, plug into this : y-y1=m(x-x1) where m is the slope, and (x1,y1) is once of your ordered pairs

OpenStudy (anonymous):

y – 11 = –3(x – 2) y – 5 = –3(x + 2) y – 5 = 3(x + 2) y – 11 = –3(x + 2)

OpenStudy (anonymous):

Just point them in a coordinate system, and look how much the y increases when the x increment by 1. Geogebra is a great program for just this

OpenStudy (amistre64):

given those points; you need something with a -11 and a +2 in it

OpenStudy (anonymous):

Those are the answer options above

OpenStudy (amistre64):

mathmater, your results are either typoed, or simply wrong :)

OpenStudy (amistre64):

then pick the one with -11 and +2 in it ....

OpenStudy (anonymous):

srry thats not point slope form

OpenStudy (anonymous):

y-5=3(x-0)

OpenStudy (amistre64):

that would be good; except they used the proper slope, and the other point :)

OpenStudy (amistre64):

consider the slope; you are higher on the left than the right so slope is negative

OpenStudy (anonymous):

its d i think